# Wasserstein Spatial Depth

### v1

@@ 4.3 Gaussian families
$\mathrm{SD}(Q;{\bf P})=\\ 1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-\boldsymbol{\mu}_{P}-\boldsymbol{A}_{Q,P}(\mathbf{x}-\boldsymbol{\mu}_{Q})}{\left(\|\boldsymbol{\mu}_{P}-\boldsymbol{\mu}_{Q}\|^{2}-{\rm Tr}\left(\boldsymbol{\Sigma}_{P}+\boldsymbol{\Sigma}_{Q}-2\left(\boldsymbol{\Sigma}^{\frac{1}{2}}_{P}\boldsymbol{\Sigma}_{Q}\boldsymbol{\Sigma}^{\frac{1}{2}}_{P}\right)^{\frac{1}{2}}\right)\right)^{\frac{1}{2}}}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
@@
which is the Euclidean spatial depth function of $\boldsymbol{\mu}_{Q}$. When ${\rm supp}(\mathbf{P})=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}}$, we obtain
$\mathrm{SD}(Q;{\bf P}):=\\ 1-\left(\int{\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\frac{\mathbf{x}-\boldsymbol{\mu}_{P_{i}}-\boldsymbol{A}_{Q,P_{i}}(\mathbf{x}-\boldsymbol{\mu}_{Q})}{\left(\|\boldsymbol{\mu}_{P_{i}}-\boldsymbol{\mu}_{Q}\|^{2}-{\rm Tr}\left(\boldsymbol{\Sigma}_{P_{i}}+\boldsymbol{\Sigma}_{Q}-2\left(\boldsymbol{\Sigma}^{\frac{1}{2}}_{P_{i}}\boldsymbol{\Sigma}_{Q}\boldsymbol{\Sigma}^{\frac{1}{2}}_{P_{i}}\right)^{\frac{1}{2}}\right)\right)^{\frac{1}{2}}}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
@@ 5.1 General properties
In this section we prove that the WSD shares the main properties of the Euclidean spatial depth, i.e., it belongs to the interval $[0,1]$, the geometric median is the deepest point, it decreases at infinite and it is transformation invariant.
@@
Then $F$ is called trivial if there exists an isometry $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ over $\mathbb{R}^{d}$ such that $F(P)=f_{\#}P$ for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$; $F$ is said to preserve shapes if for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exists an isometry $f=f_{P}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $F(P)=f_{\#}P$; and if $F$ does not preserve shapes, is said to be exotic. An example of nontrivial isometry on $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ that preserves shapes is given by the mapping $\Phi(\varphi):P\mapsto\Phi(\varphi)(P)$ where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a linear isometry and $\Phi(\varphi)(P)$ is the law of the random variable
@@
Theorems 1.1 and 1.2 in [3] proved that $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ admits exotic isometries if and only if $d=1$, which is the reason for which the invariance of the WSD holds for $d\geq 2$.
@@ 5.2 Maximality at the center
such that $Q\neq P_{i}$ for all $i=1,\dots n$ satisfies $\mathrm{SD}(Q;{\bf P})=1$.
@@ 5.3 Continuity
In this section we investigate some topological properties of the WSD. We analyze separately the function $Q\mapsto\mathrm{SD}(Q;{\bf P})$ and ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$. The following result shows that that the function $\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})\ni Q\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous.
@@
Next we show the continuity of ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$ for fixed $Q$. As in intermediate step we need to show that that for each $Q\in\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$ the function
@@
is continuous.
@@ 6 Consistent estimation
$\left(\begin{array}[]{ccc}\mathbf{X}_{1,1}&\dots&\mathbf{X}_{1,m}\\ \vdots&\vdots&\vdots\\ \mathbf{X}_{n,1}&\dots&\mathbf{X}_{n,m}\end{array}\right),$ (6)
@@ 6.1 One-stage sampling
$\mathrm{SD}(Q;{\bf P}_{n}):=1-\left(\int{\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\frac{\mathbf{x}-T_{Q,P_{i}}(\mathbf{x})}{\mathcal{W}_{2}(Q,P_{i})}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}},$
which is built upon the empirical measure ${\bf P}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}}$. The WSD is associated with the spatial distribution process
@@
pushes forward tight probability measures over $\mathcal{P}_{2}(\mathbb{R}^{d})$ to tight probabilities in $L^{2}(Q)$. Recall that a probability measure $\mu\in\mathcal{P}(\mathcal{X})$ over a separable topological space $(\mathcal{X},d)$ is said to be tight if for every $\epsilon>0$ there exists a compact (in the metric topology) set $K$ such that $\mu(K)\geq 1-\epsilon.$ A random variable is tight if its distribution is tight. Therefore, Lemma 5.5 implies that if $P\sim{\bf P}$ is tight in $\mathcal{P}_{2}(\mathbb{R}^{d})$, then $T_{Q,P}$ is tight in $L^{2}(Q)$. As a consequence, if ${\bf P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$, then $\left\{\frac{I-T_{Q,P_{i}}}{\|I-T_{Q,P_{i}}\|_{L^{2}(Q)}}\right\}_{i=1}^{n}$ is an i.d.d. sequence of tight random elements in the separable Hilbert space $L^{2}(Q)$, with finite second order moments. The strong law of large numbers and the central limit theorem in separable Hilbert spaces (cf. [24, Corollary 10.9]) yield the following result. Note that a random element $Z$ of a Hilbert space $\mathcal{H}$ is defined to be Gaussian when $h(Z)$ follows a (univariate) Gaussian distribution for all linear continuous mappings $h:\mathcal{H}\to\mathbb{R}$.
@@ 6.2 Two-stage sampling
Now we deal with the scenario where only the data (6) array is available. Recall that, in this case, the i.i.d. samples $P_{1},\dots,P_{n}$ of ${\bf P}$ are no longer observed but a sample $\{\mathbf{X}_{i,j}\}_{i,j=1}^{n,m}$ is, where $\mathbf{X}_{i,j}\sim P_{i}$ for $j\in\{1,\dots m\}$ and each $i\in\{1,\dots n\}$ We denote
${\bf P}_{n,m}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i,m}},\quad\text{with}\quad P_{i,m}:=\frac{1}{m}\sum_{j=1}^{m}\delta_{\mathbf{X}_{i,j}}\quad\text{for each}\quad i\in\{1,\dots n\}.$ (7)
@@
where $Q_{m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{X_{q,j}}$ with $X_{q,1},\ldots,X_{q,m}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$, and the convention $\frac{\boldsymbol{0}}{\boldsymbol{0}}=\boldsymbol{0}$ remains. Now we show that, as $n,m\to\infty$, ${\bf P}_{n,m}$ converges in probability in $\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ for all $p\geq 1$. We endow $\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ with the metric
@@
Let $\{{\bf P}_{n,m}\}_{n\in\mathbb{N}}$ be as in (7) where $m=m(n)$ is such that $m\to\infty$ as $n\to\infty$. Assume that ${\bf P}\in\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ for $p\geq 1$. Then
@@
A combination of Lemma 6.2, Theorem 5.6 and the continuous mapping theorem yields the following consistency result for the two-staged sample estimator.
@@
Set $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ be atomless. Let $\{{\bf P}_{n,m}\}_{n\in\mathbb{N}}$ be as in (7) where $m=m(n)$ is such that $m\to\infty$ as $n\to\infty$. Then, for every $Q\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@ 7 Comparison with other possible depth notions
In the field of nonparametric statistics, the concept of depths presents significant challenges when attempting generalization to non-Euclidean spaces, a topic that has garnered considerable attention in advanced statistical research. Within the confines of linear functional spaces, such as Banach or Hilbert spaces, the application of Euclidean methodologies remains largely successful, attributed primarily to their inherent vectorial structures. Contrastingly, the landscape becomes markedly more complex when venturing into the domain of infinite-dimensional spaces devoid of a vectorial framework.
@@ 7.1 Tukey depth
if there there exists t[0,1] such that $\text{if there there exists $t\in[0,1]$ such that }\mathcal{W}_{2}^{2}(\gamma(t),P)\leq\mathcal{W}_{2}^{2}(\gamma(t),Q)\\ \implies\left(\mathcal{W}_{2}^{2}(\gamma(0),P)\leq\mathcal{W}_{2}^{2}(\gamma(0),Q)\right)\ {\rm or}\ \left(\mathcal{W}_{2}^{2}(\gamma(1),P)\leq\mathcal{W}_{2}^{2}(\gamma(1),Q)\right).$ (8)
Recall that a constant speed geodesic in a metric space $(\mathcal{M},d)$ is a curve $\gamma:[0,1]\to\mathcal{M}$ such that $d(\gamma(t),\gamma(s))=|t-s|d(\gamma(0),\gamma(1))$ for all $s,t\in[0,1]$. In $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$, a constant speed geodesic corresponds to interpolations obtained from optimal transport plans [1, p. 158]. More precisely, any constant speed geodesic connecting two absolutely continuous distributions $P_{1}$ and $P_{2}$ is of the form $\gamma(t)=((1-t)I+tT)_{\#}P_{1}$, where $T$ is the unique optimal map pushing $P_{1}$ to $P_{2}$.
Center-outward monotonicity is widely regarded as a favorable attribute within the scope of statistical depth measures. However, it is an attribute not typically anticipated in the context of spatial depths, particularly within Euclidean spaces. Notably, neither the transport-based depth nor the lens depth exhibit this property. Moreover, for the Wasserstein halphspace depth it is not clear if the geometric condition (8), and a fortiori the center outward monotonicity, holds in general.
@@ 8 Numerical simulations
In this section, we carry out extensive numerical simulations to validate our notion of Wasserstein spatial depth and support its theoretical properties and practical utility. Specifically, we confirm the consistency of the empirical WSD, evaluate its effectiveness in outlier detection, and examine its relationship with conventional spatial depth in certain cases. Throughout this section, we use the R package transport to compute all the Wasserstein distances and optimal transport maps from data clouds. Based on the two stage sampling model in (6), the empirical WSDs are calculated via the formula below. For $Q_{m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{X_{q,j}}$ with $X_{q,1},\ldots,X_{q,m}\stackrel{{\scriptstyle\textit{i.i.d. }}}{{\sim}}Q$,
$\mathrm{SD}(Q_{m};{\bf P}_{n,m})=1-\sqrt{\frac{1}{m}\sum_{j=1}^{m}\bigg{\|}\frac{1}{n}\sum_{i\neq q}\frac{X_{q,j}-T_{Q_{m},P_{i,m}}(X_{q,j})}{\mathcal{W}_{2}(Q_{m},P_{i,m})}\bigg{\|}^{2}},$ (9)
where $Q_{m}$ could be outside (with the convention $q=n+1$) or within (with the convention $q\in\{1,\ldots,n\}$) the sampled distributions $\{P_{1,m},\ldots,P_{n,m}\}$, and where we recall the convention $/0=0$ 0. The code for all simulations is publicly available at https://github.com/YishaYao/Wasserstein-Spatial-Depth/tree/main.
Since computing the OT map between any pair of empirical distributions costs $O(m^{2})$ and once the OT map between a pair of empirical distributions is available, the corresponding Wasserstein distance immediately follows with almost zero extra cost, the computational complexity of $SD(Q_{m},\boldsymbol{P}_{n,m})$ is of order $O(nm^{2})$.

### v2

@@ 4.3 Gaussian families
$\mathrm{SD}(Q;{\bf P})=\\ 1-\left(\int{\left\|\mathbb{E}_{P\sim\mathbf{P}}\left[\frac{\mathbf{x}-\boldsymbol{\mu}_{P}-\boldsymbol{A}_{Q,P}(\mathbf{x}-\boldsymbol{\mu}_{Q})}{\left(\|\boldsymbol{\mu}_{P}-\boldsymbol{\mu}_{Q}\|^{2}+{\rm Tr}\left(\boldsymbol{\Sigma}_{P}+\boldsymbol{\Sigma}_{Q}-2\left(\boldsymbol{\Sigma}^{\frac{1}{2}}_{P}\boldsymbol{\Sigma}_{Q}\boldsymbol{\Sigma}^{\frac{1}{2}}_{P}\right)^{\frac{1}{2}}\right)\right)^{\frac{1}{2}}}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
@@
which is the Euclidean spatial depth function of $\boldsymbol{\mu}_{Q}$. When $\mathbf{P}=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}}$, we obtain
$\mathrm{SD}(Q;{\bf P}):=\\ 1-\left(\int{\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\frac{\mathbf{x}-\boldsymbol{\mu}_{P_{i}}-\boldsymbol{A}_{Q,P_{i}}(\mathbf{x}-\boldsymbol{\mu}_{Q})}{\left(\|\boldsymbol{\mu}_{P_{i}}-\boldsymbol{\mu}_{Q}\|^{2}+{\rm Tr}\left(\boldsymbol{\Sigma}_{P_{i}}+\boldsymbol{\Sigma}_{Q}-2\left(\boldsymbol{\Sigma}^{\frac{1}{2}}_{P_{i}}\boldsymbol{\Sigma}_{Q}\boldsymbol{\Sigma}^{\frac{1}{2}}_{P_{i}}\right)^{\frac{1}{2}}\right)\right)^{\frac{1}{2}}}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}}.$
@@ 5.1 General properties
In this section we prove that the WSD shares the main properties of the Euclidean spatial depth, i.e., it belongs to the interval $[0,1]$, it decreases at infinity and it is transformation invariant.
@@
Then $F$ is called trivial if there exists an isometry $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $F(P)=f_{\#}P$ for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$; $F$ is said to preserve shapes if for all $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$ there exists an isometry $f=f_{P}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $F(P)=f_{\#}P$; and if $F$ does not preserve shapes, it is said to be exotic. An example of nontrivial isometry on $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ that preserves shapes is given by the mapping $\Phi(\varphi):P\mapsto\Phi(\varphi)(P)$ where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a linear isometry and $\Phi(\varphi)(P)$ is the law of the random variable
@@
Theorems 1.1 and 1.2 in [5] prove that $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$ admits exotic isometries if and only if $d=1$, which is the reason for which the invariance of the WSD holds for $d\geq 2$.
@@ 5.2 Maximality at the center
such that $Q\neq P_{i}$ for all $i=1,\dots,n$ satisfies $\mathrm{SD}(Q;{\bf P})=1$.
@@ 5.3 Continuity
In this section we investigate some topological properties of the WSD. We analyze separately the function $Q\mapsto\mathrm{SD}(Q;{\bf P})$ and ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$. The following result shows that the function $\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})\ni Q\mapsto\mathrm{SD}(Q;{\bf P})$ is continuous.
@@
Next we show the continuity of ${\bf P}\mapsto\mathrm{SD}(Q;{\bf P})$ for fixed $Q$. As an intermediate step we need to show that for each $Q\in\mathcal{P}_{2}^{a.c.}(\mathbb{R}^{d})$ the function
@@
is continuous. Recall that $T_{Q,P}$ is the optimal transport map from $Q$ to $P$.
@@ 6 Consistent estimation
$\left(\begin{array}[]{ccc}\mathbf{X}_{1,1}&\dots&\mathbf{X}_{1,m}\\ \vdots&\vdots&\vdots\\ \mathbf{X}_{n,1}&\dots&\mathbf{X}_{n,m}\end{array}\right),$ (6.1)
@@ 6.1 One-stage sampling
$\mathrm{SD}(Q;{\bf P}_{n}):=1-\left(\int{\left\|\frac{1}{n}\sum_{i=1}^{n}\left[\frac{\mathbf{x}-T_{Q,P_{i}}(\mathbf{x})}{\mathcal{W}_{2}(Q,P_{i})}\right]\right\|^{2}}\mathrm{d}Q(\mathbf{x})\right)^{\frac{1}{2}},\quad{\bf P}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}},$ (6.2)
where $\mathbf{P}_{n}$ is the empirical counterpart to $\mathbf{P}$. The WSD is associated with the spatial distribution process
@@
pushes forward tight probability measures over $\mathcal{P}_{2}(\mathbb{R}^{d})$ to tight probabilities in $L^{2}(Q)$. Recall that a probability measure $\mu\in\mathcal{P}(\mathcal{X})$ over a separable topological space $(\mathcal{X},d)$ is said to be tight if for every $\epsilon>0$ there exists a compact (in the metric topology) set $K$ such that $\mu(K)\geq 1-\epsilon.$ A random variable is tight if its distribution is tight. Therefore, Lemma 5.5 implies that if $P\sim{\bf P}$ is tight in $\mathcal{P}_{2}(\mathbb{R}^{d})$, then $T_{Q,P}$ is tight in $L^{2}(Q)$. As a consequence, if ${\bf P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$, then $\left\{\frac{I-T_{Q,P_{i}}}{\|I-T_{Q,P_{i}}\|_{L^{2}(Q)}}\right\}_{i=1}^{n}$ is an i.i.d. sequence of tight random elements in the separable Hilbert space $L^{2}(Q)$, with finite second order moments. The strong law of large numbers and the central limit theorem in separable Hilbert spaces (cf. [33, Corollary 10.9]) yield the following result. Note that a random element $Z$ of a Hilbert space $\mathcal{H}$ is defined to be Gaussian when $h(Z)$ follows a (univariate) Gaussian distribution for all linear continuous mappings $h:\mathcal{H}\to\mathbb{R}$.
@@ 6.2 Two-stage sampling
Now we deal with the scenario where only the data array (6.1) is available. Recall that, in this case, the i.i.d. samples $P_{1},\dots,P_{n}$ of ${\bf P}$ are no longer observed but a sample $\{\mathbf{X}_{i,j}\}_{i,j=1}^{n,m}$ is, where $\mathbf{X}_{i,j}\sim P_{i}$ for $j\in\{1,\dots,m\}$ and each $i\in\{1,\dots,n\}$ We denote
${\bf P}_{n,m}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i,m}},\quad\text{with}\quad P_{i,m}:=\frac{1}{m}\sum_{j=1}^{m}\delta_{\mathbf{X}_{i,j}}\quad\text{for each}\quad i\in\{1,\dots,n\}.$ (6.3)
@@
where $Q_{m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{X_{q,j}}$ with $X_{q,1},\ldots,X_{q,m}\stackrel{{\scriptstyle\textit{i.i.d.}}}{{\sim}}Q$, and the convention $\frac{\boldsymbol{0}}{0}=\boldsymbol{0}$ remains. Now we show that, as $n,m\to\infty$, ${\bf P}_{n,m}$ converges in probability in $\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ for all $p\geq 1$. We endow $\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ with the metric
@@
Let $\{{\bf P}_{n,m}\}_{n\in\mathbb{N}}$ be as in (6.3) where $m=m(n)$ is such that $m\to\infty$ as $n\to\infty$. Assume that ${\bf P}\in\mathcal{P}(\mathcal{P}_{p}(\mathbb{R}^{d}))$ for $p\geq 1$. Then
@@
A combination of Lemma 6.2, Theorem 5.6 and the continuous mapping theorem yields the following consistency result for the two-stage sampling estimator.
@@
Set $\mathbf{P}\in\mathcal{P}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ be atomless. Let $\{{\bf P}_{n,m}\}_{n\in\mathbb{N}}$ be as in (6.3) where $m=m(n)$ is such that $m\to\infty$ as $n\to\infty$. Then, for every $Q\in\mathcal{P}^{a.c}_{2}(\mathbb{R}^{d})$,
@@
Theorems 6.1 and 6.3 state that the empirical WSD converges to the population version asymptotically. Given enough sample size, the empirical WSD is informative of the truth and has practical values. The simulation results in Section 8.1 also verify the above theorems.
@@ 7 Comparison with other possible depth notions
In the field of nonparametric statistics, the concept of depth presents significant challenges when attempting generalization to non-Euclidean spaces, a topic that has garnered considerable attention in advanced statistical research. Within the confines of linear functional spaces, such as Banach or Hilbert spaces, the application of Euclidean methodologies remains largely successful, attributed primarily to their inherent vectorial structures. Contrastingly, the landscape becomes markedly more complex when venturing into the domain of infinite-dimensional spaces devoid of a vectorial framework.
@@ 7.1 Tukey depth
[…]
Recall (Section 3.2) that a constant speed geodesic in a metric space $(\mathcal{M},d)$ is a curve $\gamma:[0,1]\to\mathcal{M}$ such that $d(\gamma(t),\gamma(s))=|t-s|d(\gamma(0),\gamma(1))$ for all $s,t\in[0,1]$. In $(\mathcal{P}_{2}(\mathbb{R}^{d}),\mathcal{W}_{2})$, a constant speed geodesic corresponds to interpolations obtained from optimal transport plans [1, p. 158]. More precisely, any constant speed geodesic connecting two absolutely continuous distributions $P_{1}$ and $P_{2}$ is of the form $\gamma(t)=((1-t)I+tT_{P_{1},P_{2}})_{\#}P_{1}$, where $T_{P_{1},P_{2}}$ is the unique optimal map pushing $P_{1}$ to $P_{2}$ (see also Section 3.3).
Center-outward monotonicity is widely regarded as a favorable attribute within the scope of statistical depth measures. However, it is an attribute not typically anticipated in the context of spatial depths, particularly within Euclidean spaces. Notably, neither the transport-based depth nor the lens depth exhibit this property. Moreover, for the Wasserstein halphspace depth it is not clear if the geometric condition (7.1), and a fortiori the center-outward monotonicity, holds in general.
@@ 8 Numerical simulations
In this section, we carry out extensive numerical simulations to validate our notion of Wasserstein spatial depth and support its theoretical properties and practical utility. Specifically, we confirm the consistency of the empirical WSD, examine its relationship with conventional spatial depth in certain cases, evaluate its effectiveness in outlier detection and show its benefit compared to applying functional depths to distributions. Throughout this section, we use the R package transport to compute all the Wasserstein distances and optimal transport maps from data clouds. Based on the two stage sampling model in (6.1), the empirical WSDs are calculated via the formula below. For $Q_{m}=\frac{1}{m}\sum_{j=1}^{m}\delta_{X_{q,j}}$ with $X_{q,1},\ldots,X_{q,m}\stackrel{{\scriptstyle\textit{i.i.d. }}}{{\sim}}Q$,
$\mathrm{SD}(Q_{m};{\bf P}_{n,m})=1-\sqrt{\frac{1}{m}\sum_{j=1}^{m}\bigg{\|}\frac{1}{n_{q}}\sum_{i\neq q}\frac{X_{q,j}-T_{Q_{m},P_{i,m}}(X_{q,j})}{\mathcal{W}_{2}(Q_{m},P_{i,m})}\bigg{\|}^{2}},$ (8.1)
[…]
Since computing the optimal transport map between any pair of empirical distributions costs $O(m^{2})$ [49], and once the optimal transport map between a pair of empirical distributions is available, the corresponding Wasserstein distance immediately follows with almost zero extra cost, the computational complexity of $\mathrm{SD}(Q_{m},\boldsymbol{P}_{n,m})$ is of order $O(nm^{2})$.